# Latent Beam Diffusion Models for Decoding Image Sequences

### v1

@@ Appendix A Impact of Latent Selection on Beam Search Performance
Latent selection plays a crucial role in determining the quality and coherence of image sequences generated through beam search. By analyzing how different latent groups contribute to generation performance, we can better understand their impact on semantic alignment and visual coherence. To investigate this, we evaluate 15 tasks, each consisting of four steps, allowing us to systematically assess the effects of latent selection across a diverse set of sequences. Using both automatic evaluation and human annotations, we analyze different latent configurations to determine which groups contribute most to generating coherent and semantically meaningful sequences.
A.1 Impact of Latent Groups on Performance
@@
Following the initial analysis of beam search configurations (Sec. 5.2), we further investigated the impact of using different latent configurations on the generated sequences. While latents 0, 1, 2 and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance. For these latents, the beam search configurations yield very limited selection rates, indicating that they are rarely selected in both human and Gemini annotations. Specifically, for human annotations evaluated across 6 tasks, the only meaningful result is from the configuration with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%. Other configurations show no selection rates.
@@
Similarly, for Gemini annotations evaluated on the same 6 tasks, the configurations do not yield any notable results, with no selection rates. When the evaluation is extended to 35 tasks, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%. Figure 8 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [32, 4].
@@
Given the negative impact of higher latents on performance, as observed in human annotations, we decided to investigate the underlying reasons for this trend. Specifically, we explored how latents affect text-image alignment using the DAAM (Diffusion Attention Attribution Map) framework [44]. DAAM helps interpret text-to-image diffusion models by analyzing the cross-attention maps within the denoising process. It generates heat maps that highlight how much each token in the prompt influences different parts of the generated image, offering a visual representation of text-image alignment. By applying DAAM, we aimed to determine whether higher latents distort or weaken the contribution of key tokens during generation.
We then used DAAM to analyze the cross-attention between the input tokens and the generated image. DAAM enables us to track how much each remaining token influences specific regions of the image by calculating the attention patterns in the model. The resulting heat maps visualize this influence, with the intensity of the heat maps indicating the level of contribution each token has on different parts of the generated image. This approach helps us understand how well the model attends to key parts of the prompt and ensures that the most meaningful tokens are appropriately reflected in the image.
To quantify text alignment, we calculated the average heat map maximum intensity for all remaining tokens, providing a numerical measure of how well the model preserved the key textual elements in the generated image. Then, for each sequence of generated images, we computed the average alignment score across all images in the sequence. While this serves as an objective metric for evaluating text adherence, it remains a partial measure, as it focuses solely on text alignment without accounting for other important aspects of image quality, such as coherence, realism, or artistic fidelity. Figure 9 shows the average impact of different latent groups on text alignment. The results reveal a clear pattern, latents from earlier iterations contribute more effectively to aligning the text with the generated image, while text alignment weakens as we move to higher latent ranges. Specifically, latents in the 0–3 range showed the strongest alignment, with an average score of 1.544, followed by the 4–7 range (1.523) and the 8–11 range (1.435). In contrast, latents from the highest index range (16–19) had the weakest influence on alignment, with a significant drop in average scores. This suggests that earlier-stage latents play a crucial role in preserving text fidelity, whereas later-stage latents may dilute or distort the contribution of key tokens.
@@
While later latents are effective at preserving visual consistency from previous steps, they are less capable of incorporating new details from the input prompts. For example, later latent groups (4-7) maintain stability in visual elements throughout the process, preserving previously established features. However, this comes at the expense of adapting to new or changing textual instructions. In contrast, early latent groups (0-3) demonstrate strong alignment between the text and the generated image, with consistently high heat map intensity across all steps. This highlights their role in grounding the image in the input text. For later latent groups (4-7), however, we observe a drop in heat map intensity, particularly by step 3, suggesting the model shifts focus away from strict textual adherence and toward refining finer visual details.
A.2 Latent Selection on Image Sequence Generation
@@
Following our evaluation of the impact of latent groups on performance, we now examine the influence of individual latents within the groups that exhibit better text alignment (0-3). As shown in Figure 10, Latent 3 is used most frequently, accounting for 44.4% of selections, followed by Latent 0 at 20%. In contrast, Latents 1 and 2 are used less often, with Latent 1 at 11.1% and Latent 2 at 8.9%. This uneven distribution suggests that Latents 0 and 3 are particularly effective at capturing key features at critical points in the generation process, which beam search exploits to enhance sequence coherence. The ”Rand seed,” which introduces variability into the process, accounts for 15.6% of selections. This highlights beam search’s flexibility, allowing it to explore a range of possible outputs while preserving the overall structure of the image sequence.
In general, these findings reinforce the importance of latent selection strategies in optimizing image generation, suggesting that models could benefit from prioritizing the most impactful latents to achieve both coherent visual sequences and stronger text-image alignment.
@@
Figure 11 illustrates the impact of different latent search methods on text alignment. The heat maps reveal how attention patterns shift based on the chosen decoding strategy, highlighting the differences in how each approach preserves or redistributes focus on textual cues throughout the image generation process. As shown, particularly in step 3, BeamDiffusion demonstrates superior text alignment compared to the other methods. Not only does it maintain a stronger connection to the input text, it also ensures better overall sequence coherence. This improved alignment and consistency across generations are key advantages of BeamDiffusion.
These variations provide further insight into how latent search methods can influence both the coherence of the generated sequence and the alignment between the input text and visual output.
Appendix B Minimizing Decoding Risk
@@
The sequence of scenes $S=\{s_{1},s_{2},\dots,s_{l}\}$ may have one scene in the middle of the sequence that should be independent. To address this, for steps $j>1$, we use a combination of random seeds and past latent vectors to maintain a balance between diversity and coherence, which is particularly important for such scenarios.
Appendix C Annotation Tasks
Human annotations.
  • Steps Back: This refers to how many previous steps the model uses from its latent representations to explore new candidate outputs. We evaluated three different settings for the number of steps back: 2, 3, and 4, which correspond to using latents from the last two, three, or four images, respectively.
  • Number of Latents: We tested two sets of latent configurations {1, 3, 5, 7} and {0, 1, 2, 3} to explore how varying levels of latent detail impact the quality of the generated sequences.
  • Beam Width: We explored four different beam widths {2, 3, 4, 6} which determine how many candidate sequences are retained at each step during the search process.
Evaluating all of these at once proved to be exhausting and difficult to manage. To simplify the evaluation process, we adopted a ”battle-style” annotation method. In this approach, 5 annotators compare two configurations at a time and select the one they find has better overall coherence. If neither configuration stands out, annotators can select ”both good” or ”both bad” (Figure 12), in which case the computationally cheaper option advances to the next round. This process is repeated until we identify the best configuration. This method focuses on direct comparisons, making it easier to pinpoint the most effective beam search settings while also considering computational efficiency.
Figure 12: Example of the human annotation battle round for selecting the best beam search configuration.
To evaluate all models, 10 annotators are presented with four generated sequences side by side: one from the top-performing beam search configuration (selected during the initial annotations) and three others from alternative decoding methods, such as greedy search, nucleus sampling, and GILL. Figure 13 illustrates an example of this comparison process, where the annotators assess the sequences to determine the most effective approach.
Figure 13: Example of a human annotation to choose the best decoding method.
Gemini evaluation.
Figure 14 shows the prompt used in Gemini to annotate the best configuration and the best method. This prompt was used to guide the evaluation of image sequences based on visual and semantic consistency.
Gemini Prompt: ”Evaluate two image sequences using the following criteria: Visual Consistency: Objects, ingredients, backgrounds, colors and styles should remain consistent across frames. Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements. Semantic Consistency: Images must logically follow the described steps, clearly and accurately. Steps of the process: {steps} After evaluating all sequences, select the one with the strongest overall consistency (visual and semantic). Pay close attention to even minimal differences, as the sequences are very similar. If both sequences are equally good or bad, choose the one that is slightly better. Choose exactly one of the following options: - ”First sequence” - ”Second sequence” ”
Appendix D Datasets
Recipes and DIYs.
VIST.
Appendix E Challenges
@@
Captions.
One of the challenges encountered in this process is generating captions that accurately describe the steps involved in the sequence. Often, a single step may consist of multiple sub-steps, each contributing to the overall generation. This makes it difficult to write a single concise caption that captures the full scope of each step. The challenge lies in ensuring that the caption not only reflects the specific action or transformation occurring at each stage but also conveys the progression and complexity of the entire sequence. Additionally, if we divide the steps into smaller sub-steps to better describe the process, the sequence becomes very extensive, making it even harder to maintain clarity and readability while ensuring comprehensive coverage of the entire generation process.
Domains.
@@
Complexity of the generation.
The Beam Diffusion model generates multiple candidate images at each step by exploring a range of latents. We analyze the complexity of the model for a steps back of $n$ steps and $k$ latents explored. In the initial steps, the complexity is high because pruning begins only after the second step. For each beam, $k\times n$ latents are explored, and the search space grows exponentially, resulting in significant computational cost in the early stages. During the first two steps, the complexity grows with the number of candidates and latents, proportional to $O((k\times n)^{j})$ for $j\leq 2$. Specifically, for $j=1$, the complexity is proportional to the number of explored random seeds. After the second step, pruning reduces the number of candidates to the most promising $w$ beams, decreasing the complexity to $O(k\times n\times w)$.
To further reduce complexity, a heuristic based on text relevance can be used to predict the top $T$ most relevant past images. By restricting the search space to the latents of these top $T$ candidates, the model improves efficiency by prioritizing contextually relevant images, particularly in the early steps. While the initial three steps remain computationally demanding due to the expansive search space, pruning combined with this heuristic can significantly lower the computational cost by refining the candidate set.
@@ Appendix F Application Examples
Images and other visual aids are essential tools that significantly improve comprehension in various fields. Whether while breaking down complex tasks, illustrating a sequence of events, or clarifying intricate concepts, visual representations provide clear, step-by-step guidance that makes the material easier to understand [26]. From technical procedures and instructional design to storytelling and educational content, the use of multiple-scene images allows the reader to grasp information more deeply and intuitively, transforming even the most complex or dense topics into something more accessible [39, 11]. However, generating a sequence of images that not only aligns with each textual instruction but also maintains overall coherence is a major challenge  [25].
Figure 15: Visualization of cross-attention in latent diffusion across different sets of latents. The heat maps reveal how attention patterns evolve at various stages, highlighting the alignment between text prompts and distinct latent representations.
@@ F.1 Beam Search Tree
In Figure 16(b), we observe that the generation of the selected image at step 4 behaves similarly to an isolated diffusion process. However, it is important to note that this process is influenced by multiple prompts. Specifically, the image generation at step 4 is conditioned on the prompts from steps 1, 3, and 4. This demonstrates how the latent search process incorporates context from various stages, showing that even though each generation step seems independent, it is shaped by previous steps, enabling a more coherent and consistent generation of the final image.
@@ F.2 Sequence Examples
To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 20-28. These figures show case how each method performs in terms of maintaining sequence coherence throughout sequential image generation.
@@
In our overall qualitative analysis, we observe that both Greedy/CoSeDlen=1 and Nucleus Sampling methods are generally good at following the textual input across steps, as shown in Figures 20, 20, 24, and 24. They align well with the step instructions but often struggle with visual consistency, as the generated images show noticeable shifts or inconsistencies between steps. This lack of visual cohesion detracts from the overall coherence of the sequence, even though they adhere reasonably well to the textual prompt. In contrast, GILL exhibits a different behavior. While it maintains high visual consistency, with all images in the sequence being visually similar, it fails to accurately follow the step-by-step textual instructions. GILL tends to generate images that lack the necessary variability to incorporate new information from each step, resulting in sequences that are visually cohesive but contextually incorrect or incomplete.
@@
Finally, BeamDiffusion strikes a strong balance between text adherence and visual consistency. As shown in Figures 20, 20, 24, and 24, it generally produces the most coherent sequences compared to other methods. In most cases, BeamDiffusion effectively follows the textual instructions while maintaining visual consistency across the steps. This combination of textual alignment and visual stability leads to more coherent and contextually accurate image sequences. While not without its limitations, BeamDiffusion demonstrates a more reliable progression throughout the image generation process than the other approaches. However, it is important to note that all methods face challenges in maintaining coherence within the DIY domain, as seen in Figures 28, 28, and 28. The complexity and diversity of DIY tasks present unique difficulties for the models, making it harder to both accurately follow textual instructions and ensure consistency throughout the sequence. While the models perform well in other domains, the intricacies of the DIY domain result in some inconsistencies, particularly when capturing step changes.
@@
Figure 17: Example of the Recipes domain task ”Vegan Basil Pesto and Sun-Dried Tomato Pinwheels”, with a sequence of 4 steps.
@@
Figure 18: Example of the Recipes domain task ”Tangy lemon sheet cake”, with a sequence of 4 steps.
@@
Figure 19: Example of the Recipes domain task ”Zucchini, Pecorino & Mint Salad”, with a sequence of 4 steps.
@@
Figure 20: Example of the Recipes domain task ”Bone Broth”, with a sequence of 3 steps.
@@
Figure 21: Example of the out-of-domain VIST story ”Klamath Project Construction”, with a sequence of 5 steps.
@@
Figure 22: Example of the out-of-domain VIST story ”Yosemite Weekend”, with a sequence of 5 steps.
@@
Figure 23: Example of the out-of-domain VIST story ”Fireworks - 4th July in Washington, DC”, with a sequence of 5 steps.
@@
Figure 24: Example of the out-of-domain VIST story ”Independence Day Parade”, with a sequence of 5 steps.
@@
Figure 25: Example of the out-of-domain DIY task ”How to Clean Silk Rugs”, with a sequence of 2 steps.
@@
Figure 26: Example of the out-of-domain DIY task ”How to Polish Wood”, with a sequence of 3 steps.
@@
Figure 27: Example of the out-of-domain DIY task ”How to Preserve Basil”, with a sequence of 6 steps.
@@
Figure 28: Example of the out-of-domain DIY task ”How to Clean Natural Stone”, with a sequence of 5 steps.

### v2

@@ Appendix C Impact of Latent Selection on Beam Search Performance
Latent selection plays a crucial role in determining the quality and coherence of image sequences generated through beam search. By analyzing how different latent groups contribute to generation performance, we can better understand their impact on semantic alignment and visual coherence.
This aligns with recent work on interpreting diffusion models. For example, Pont-Tuset et al. [30] ground textual descriptions in image regions, while Tang et al. [46] and Dewan et al. [5] explore how textual prompts influence generation through attention analysis and partial information decomposition. In contrast, our method contributes to interpretability by examining how latent space traversal, guided by beam search, affects the generation of coherent, semantically aligned image sequences.
To systematically investigate this, we evaluate several sequences, each consisting of four steps, allowing us to assess the effects of latent selection across a diverse set of sequences. Using both automatic evaluation and human annotations, we analyze different latent configurations to determine which groups contribute most to generating coherent and semantically meaningful sequences.
C.1 Impact of Latent Groups on Performance
@@
Following the initial analysis of beam search configurations (Sec. 5.3), where we used latent indexes 0, 1, 2, and 3, we further investigated the impact of using different latent configurations on the generated sequences. While latents 0, 1, 2, and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance. In particular, beam search configurations for these underperforming latents result in very limited selection rates, indicating they are rarely chosen in both human and Gemini annotations. For human annotations across 6 tasks, the only configuration that yielded any selection was the one with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%. All other configurations resulted in 0% selection. In contrast, none of the configurations showed any non-zero selection rate under Gemini annotations, reinforcing the lack of preference for these latents.
@@
Similarly, for Gemini annotations evaluated on the same 6 sequences, the configurations do not yield any notable results, with no selection rates. When the evaluation is extended to 35 sequences, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%. Figure 8 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [33, 2].
@@
Given the observed negative impact of higher latents on performance in human annotations, we investigated how latents influence text-image alignment using the DAAM (Diffusion Attention Attribution Map) framework [46]. DAAM interprets text-to-image diffusion models by analyzing cross-attention maps during the denoising process, generating heat maps that reveal how strongly each prompt token affects different regions of the output image.
By applying DAAM, we examined whether higher latents distort or weaken the influence of key tokens during generation. The resulting heat maps allowed us to visually assess token-level contributions, helping us understand whether the model continues to attend to the most semantically important parts of the prompt under varying latent configurations..
To quantify text alignment, we calculated the average heat map maximum intensity for all tokens, providing a numerical measure of how well the model preserved the key textual elements in the generated image. Then, for each sequence of generated images, we computed the average alignment score across all images in the sequence. While this serves as an objective metric for evaluating text adherence, it remains a partial measure, as it focuses solely on text alignment without accounting for other important aspects of image quality, such as coherence, realism, or artistic fidelity. Figure 9 shows the average impact of different latent groups on text alignment. The results reveal a clear pattern, latents from earlier iterations contribute more effectively to aligning the text with the generated image, while text alignment weakens as we move to higher latent ranges. Specifically, latents in the 0–3 range showed the strongest alignment, with an average score of 1.544, followed by the 4–7 range (1.523). In contrast, latents from the highest index range (16–19) had the weakest influence on alignment, with a significant drop in average scores. This suggests that earlier-stage latents play a crucial role in preserving text fidelity, whereas later-stage latents may dilute or distort the contribution of key tokens.
@@
While later latents are effective at preserving visual consistency from previous steps, they are less capable of incorporating new details from the input prompts. For example, later latent groups (16-19) maintain stability in visual elements throughout the process, preserving previously established features. However, this comes at the expense of adapting to new or changing textual instructions. In contrast, early latent groups (0-3) demonstrate strong alignment between the text and the generated image, with consistently high heat map intensity across all steps. This highlights their role in grounding the image in the input text. For later latent groups (4-7), however, we observe a drop in heat map intensity, particularly by step 3, suggesting the model shifts focus away from strict textual adherence and toward refining finer visual details.
C.2 Latent Selection on Image Sequence Generation
@@
Following our evaluation of the impact of latent groups on performance, we now examine the role of individual latents within the higher-performing group (Latents 0-3). As shown in Figure 10, Latent 3 is selected most frequently (44.4%), followed by Latent 0 (20%), while Latents 1 and 2 are used less often (11.1% and 8.9%, respectively). This uneven distribution suggests that Latents 0 and 3 contribute disproportionately to the generation of sequences with better text alignment. One possible explanation is that these latents encode more semantically relevant or visually discriminative features that align well with the text prompts. Their frequent selection by beam search indicates that these latents are particularly effective at resolving ambiguities or reinforcing important visual cues during sequence generation. The lower usage of Latents 1 and 2 may reflect their tendency to encode less salient or redundant information, which contributes less to text-conditioned coherence. Interestingly, the Rand seed, which injects stochasticity into the generation process, accounts for 15.6% of selections. This suggests that while beam search primarily exploits strong latents like 0 and 3, it also values diversity. The inclusion of randomized elements allows the search to escape local optima and sample from alternative yet plausible paths, enhancing the robustness of the final sequence.
@@
Figure 11 illustrates how different latent search methods affect text-image alignment. The heat maps show how attention patterns shift depending on the decoding strategy, revealing variations in how well each method preserves focus on key textual cues. Notably, in step 3, BeamDiffusion exhibits superior alignment, maintaining a stronger connection to the prompt and ensuring greater sequence coherence compared to other methods.
While Beam Diffusion generally achieves better visual consistency and stronger text adherence in earlier steps, it slightly underperforms in step 5, failing to capture the "fireworks" mentioned in the prompt. This highlights how even methods with strong overall performance can exhibit localized weaknesses, underscoring the importance of analyzing alignment step by step.
Overall, these findings highlight the significance of latent selection strategies in optimizing image generation. The results indicate that prioritizing the most significant latents could enhance the coherence of visual sequences and improve text alignment. These observations underline the influence of latent search strategies on the semantic accuracy and coherence of generated outputs.
C.3 Effect of Cumulative Latent Access
To investigate how access to latents from previous steps influences the quality of generation, we conducted an ablation study in which we constrain the set of available latents throughout the generation of the sequence. Specifically, we evaluate performance when the model is limited to latents from only the first step, then from the first two steps, and so on, up to including latents from all previous steps (i.e., full cumulative access).
Figure 12: Visual comparison of sequences generated with cumulative latent access. Each row shows a sequence generated using latents from only the first step, the first two steps, and so on, up to all previous steps. The bottom row shows the result from BeamDiffusion, which uses latents from only the two preceding steps.
As shown in Figure 12, limiting the model to latents from only the first step results in weak text-image alignment and incoherent visual transitions. Gradually expanding the latent set to include more steps improves sequence quality, particularly when moving from two to three steps. However, we observe that including latents from all previous steps does not yield further improvements in coherence, suggesting diminishing returns with increasing temporal access.
In contrast, BeamDiffusion uses latents from only the two most recent steps and achieves results that are comparable, or even superior, to the full cumulative case. This indicates that a narrow temporal window is sufficient to maintain strong alignment and visual consistency. By avoiding the computational cost of considering all previous steps, BeamDiffusion finds a favorable trade-off between quality and efficiency.
Overall, this experiment highlights that cumulative latent access enhances generation quality up to a point, but full historical memory is unnecessary. Strategies such as BeamDiffusion, which take advantage of a limited but informative temporal context, can produce high-quality sequences while significantly reducing computational overhead. These findings also underscore the importance of the non-linear sequence denoising strategy introduced in Section 5.4, which enables more targeted and effective reuse of past latents without the need for full sequence access.
Appendix D Minimizing Decoding Risk
@@
The sequence of scenes $S=\{s_{1},s_{2},\dots,s_{L}\}$ may have one scene in the middle of the sequence that should be independent. To address this, for steps $s_{j>1}$, we use a combination of random seeds and past latent vectors to maintain a balance between diversity and coherence, which is particularly important for such scenarios.
Appendix E Broader Impacts and Challenges
@@
E.1 Impacts
This work proposes a method to improve generative pretraining by incorporating subject-aware mutual information maximization, aiming to produce more contextually aligned and coherent sequences of images. The ability to generate visually consistent image sequences has positive applications in areas such as animation, visual storytelling, educational media, and accessibility tools (e.g., for generating illustrative content from text). However, as with many generative models, there are potential negative societal impacts. Enhanced coherence and subject fidelity could be misused to create misleading or deceptive visual content, such as more convincing visual disinformation, propaganda, or synthetic media (e.g., deepfakes).
E.2 Contextual Scene Descriptions
One of the challenges encountered concerns the generation of captions that accurately describe the steps involved in the sequence. Often, a single step may consist of multiple sub-steps, each contributing to the overall generation. This makes it difficult to write a single concise caption that captures the full scope of each step. The challenge lies in ensuring that the caption not only reflects the specific action or transformation occurring at each step but also conveys the progression and complexity of the entire sequence. Additionally, if we divide the steps into smaller sub-steps to better describe the process, the sequence becomes very extensive, making it even harder to maintain clarity and readability while ensuring comprehensive coverage of the entire generation process.
To address this challenge, we leverage Gemini [34], specifically the Gemini 1.5 Flash model, to refine step descriptions into visually detailed prompts while considering sequence context. Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
$\small c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (7)
Following this approach we can improve image generation prompts by maintaining consistency and preserving the dependencies between sequential steps.
E.3 Domains.
@@ E.2 Contextual Scene Descriptions
E.4 Complexity of the generation.
The BeamDiffusion model generates multiple candidate images at each step by exploring a range of latents. We analyze the complexity of the model for a steps back of $m$ steps and $N$ latents explored. In the initial steps, the complexity is high because pruning begins only after the second step. For each beam, $N\cdot m$ latents are explored, and the search space grows exponentially, resulting in significant computational cost in the early stages. During the first two steps, the complexity grows with the number of candidates and latents, proportional to $\mathcal{O}(N\cdot m)^{j})$ for $j\leq 2$. Specifically, for $j=1$, the complexity is proportional to the number of explored random seeds. After the second step, pruning reduces the number of candidates to the most promising $w$ beams, decreasing the complexity to $\mathcal{O}(N\cdot m\cdot w)$.
To further reduce complexity, a heuristic based on text relevance can be used to predict the top $P$ most relevant past images. By restricting the search space to the latents of these top $P$ candidates, the model improves efficiency by prioritizing contextually relevant images, particularly in the early steps. While the initial three steps remain computationally demanding due to the expansive search space, pruning combined with this heuristic can significantly lower the computational cost by refining the candidate set.
@@ Appendix F Application Examples
This section presents additional visual examples and qualitative results that complement the main paper. Images and other visual aids are essential tools that significantly improve comprehension across a variety of domains. Whether breaking down complex tasks, illustrating a sequence of events, or clarifying intricate concepts, visual representations provide clear, step-by-step guidance that enhances understanding [24]. From technical procedures and instructional design to storytelling and educational content, multi-scene image sequences help readers grasp information more intuitively, transforming even the most complex or dense topics into something more accessible. However, generating a sequence of images that not only aligns with each textual instruction but also maintains overall coherence remains a major challenge [23]. In this appendix, we include expanded examples to further illustrate how our model addresses this issue. These include additional visualizations of the beam search process and extended qualitative comparisons across domains such as recipes, visual storytelling, and DIY tasks. Together, these examples highlight the strengths and limitations of current approaches in achieving textual alignment and visual consistency in sequential image generation.
Figure 13: Visualization of cross-attention in latent diffusion across different sets of latents. The heat maps reveal how attention patterns evolve at various sequence steps, highlighting the alignment between text prompts and distinct latent representations.
@@ F.1 Beam Search Tree
In Figure 14(b), we observe that the generation of the selected image at step 4 behaves similarly to an isolated diffusion process. However, it is important to note that this process is influenced by multiple prompts. Specifically, the image generation at step 4 is conditioned on the prompts from steps 1, 3, and 4. This demonstrates how the latent search process incorporates context from various steps, showing that even though each generation step seems independent, it is shaped by previous steps, enabling a more coherent and consistent generation of the final image.
@@ F.2 Sequence Examples
To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 18-26. These figures show case how each method performs in terms of maintaining sequence coherence throughout sequential image generation.
@@
In our overall qualitative analysis, we observe that both Greedy/CoSeDlen=1 and Nucleus Sampling methods are generally good at following the textual input across steps, as shown in Figures 18, 18, 22, and 22. They align well with the step instructions but often struggle with visual consistency, as the generated images show noticeable shifts or inconsistencies between steps. This lack of visual cohesion detracts from the overall coherence of the sequence, even though they adhere reasonably well to the textual prompt. In contrast, GILL and StackedDiffusion exhibit a different behavior. While both maintain high visual consistency, generating sequences in which the images are visually similar, they struggle to accurately follow the step-by-step textual instructions. GILL, in particular, tends to produce images that lack the necessary variability to incorporate new information from each step, resulting in sequences that are visually cohesive but contextually incorrect or incomplete. StackedDiffusion performs slightly better than GILL in this regard, capturing some of the step-wise changes, but still often fails to fully reflect the evolving textual input across the sequence.
@@
Finally, BeamDiffusion strikes a strong balance between text adherence and visual consistency. As shown in Figures 18, 18, 22, and 22, it generally produces the most coherent sequences compared to other methods. In most cases, BeamDiffusion effectively follows the textual instructions while maintaining visual consistency across the steps. This combination of textual alignment and visual stability leads to more coherent and contextually accurate image sequences. While not without its limitations, BeamDiffusion demonstrates a more reliable progression throughout the image generation process than the other approaches. However, it is important to note that all methods face challenges in maintaining coherence within the DIY domain, as seen in Figures 26, 26, and 26. The complexity and diversity of DIY tasks present unique difficulties for the models, making it harder to both accurately follow textual instructions and ensure consistency throughout the sequence. While the models perform well in other domains, the intricacies of the DIY domain result in some inconsistencies, particularly when capturing step changes.
@@
Figure 15: Example of the Recipes domain task "Vegan Basil Pesto and Sun-Dried Tomato Pinwheels", with a sequence of 4 steps.
@@
Figure 16: Example of the Recipes domain task "Tangy lemon sheet cake", with a sequence of 4 steps.
@@
Figure 17: Example of the Recipes domain task "Zucchini, Pecorino & Mint Salad", with a sequence of 4 steps.
@@
Figure 18: Example of the Recipes domain task "Bone Broth", with a sequence of 3 steps.
@@
Figure 19: Example of the out-of-domain VIST story "Klamath Project Construction", with a sequence of 5 steps.
@@
Figure 20: Example of the out-of-domain VIST story "Yosemite Weekend", with a sequence of 5 steps.
@@
Figure 21: Example of the out-of-domain VIST story "Fireworks - 4th July in Washington, DC", with a sequence of 5 steps.
@@
Figure 22: Example of the out-of-domain VIST story "Independence Day Parade", with a sequence of 5 steps.
@@
Figure 23: Example of the out-of-domain DIY task "How to Clean Silk Rugs", with a sequence of 2 steps.
@@
Figure 24: Example of the out-of-domain DIY task "How to Polish Wood", with a sequence of 3 steps.
@@
Figure 25: Example of the out-of-domain DIY task "How to Preserve Basil", with a sequence of 6 steps.
@@
Figure 26: Example of the out-of-domain DIY task "How to Clean Natural Stone", with a sequence of 5 steps.